OpenStudy (anonymous):

can someone please help me with this implicit differentiation problem? I have to find dy/dx or y' xy^3 + 4x^2y^2-5xy=3x+2y

OpenStudy (amistre64):

oy, thats just designed to give you tons of practice now isnt it ....

OpenStudy (amistre64):

its just alot of product rules and power rules .... just have to keep track of it is all

OpenStudy (anonymous):

haha yeah, i am horrrrible at them :/

OpenStudy (amistre64):

xy^3, x'y^3+3xy^2y' +4x^2y^2, 8xx' y^2+8x^2y y' -5xy, -5x'y -5xy' = 3x, 3x' +2y, 2y' since x' = dx/dx = 1, all those x' parts can be ignored

OpenStudy (amistre64):

\[y^3+3xy^2y'+8xy^2+8x^2y y'-5y -5xy' = 3+2y'\] gather your y's together and algebra it out

OpenStudy (amistre64):

* your (y')s for clarity

OpenStudy (anonymous):

haha i thinkkkk i solved it? does it equal (3+5y-8xy^2 -3y)/ (-2-5+8x^2y+3xy^2) ??

OpenStudy (amistre64):

i dont see the y^3 in that setup

OpenStudy (amistre64):

\[y^3+\color{red}{3xy^2y'}+8xy^2+\color{red}{8x^2y y'}-5y -\color{red}{5xy'} = 3+\color{red}{2y'}\] \[\color{red}{3xy^2y'}+\color{red}{8x^2y y'}-\color{red}{5xy'}-\color{red}{2y'} = 3-y^3-8xy^2+5y\] \[y'(\color{red}{3xy^2}+\color{red}{8x^2y}-\color{red}{5x}-\color{red}{2)} = 3-y^3-8xy^2+5y\] \[y'=\frac{3-y^3-8xy^2+5y}{\color{red}{3xy^2}+\color{red}{8x^2y}-\color{red}{5x}-\color{red}{2} }\]

OpenStudy (amistre64):

its just alot to keep track of ... :)

OpenStudy (anonymous):

i finally got that but i got a -y^3 not -y^2 in the numerator?

OpenStudy (amistre64):

xy^3 derives to y^3 + 3xy^2 y' which is where the y^3 comes from

OpenStudy (anonymous):

ohhh okay so the final answer is 3+ 5y-8xy^2-y^3)/(3xy^2+8x^2y-5x-3)?

OpenStudy (amistre64):

3+ 5y -8xy^2 -y^3 tops good ------------------- 3xy^2 +8x^2y -5x -3 -2, not -3 on the end there

OpenStudy (anonymous):

youuuu are a lifesaverrrrrrrr. (:

OpenStudy (anonymous):

could you help me with onee more implicit differentiation problem?

OpenStudy (amistre64):

i dunno, i had to dig deep just to get thru that one :)

OpenStudy (anonymous):

hahaha